# The first catalogue of candidate white dwarf-main sequence binaries in open star clusters: A new window into common envelope evolution

### v1

@@ Abstract
Close binary systems are the progenitors to both Type Ia supernovae and the compact object mergers that can be detected via gravitational waves. To achieve a binary with a small radial separation, it is believed that the system likely undergoes common envelope (CE) evolution. Despite its importance, CE evolution may be one of the largest uncertainties in binary evolution due to a combination of computational challenges and a lack of observed benchmarks where both the post-CE and pre-CE conditions are known. Identifying post-CE systems in star clusters can partially circumvent this second issue by providing an independent age constraint on the system. For the first time, we conduct a systematic search for white dwarf (WD) and main-sequence (MS) binary systems in 299 Milky Way open star clusters. Coupling Gaia DR3 photometry and kinematics with multi-band photometry from Pan-STARRS1 and 2MASS, we apply a machine learning based approach and find 52 high-probability candidates in 38 open clusters. For a subset of our systems, we present follow-up spectroscopy from the Gemini and Lick Observatories and archival light curves from TESS, Kepler/K2 and the Zwicky Transient Facility. Examples of M-dwarfs with hot companions are spectroscopically observed, along with regular system variability. While the kinematics of our candidates are consistent with their host clusters, some systems have spatial positions offset relative to their hosts, potentially indicative of natal kicks. Ultimately, this catalogue is a first step to obtaining a set of observational benchmarks to better link post-CE systems to their pre-CE progenitors.
@@
journal: ApJfacilities: Gaia, Gemini (GMOS), Kepler/K2, Lick (Kast), Pan-STARRS, TESS, ZTF, 2MASSsoftware: astropy (Astropy Collaboration et al., 2022), lightkurve (Lightkurve Collaboration et al., 2018), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010), pyraf (Science Software Branch at STScI, 2012), scikit-learn (Pedregosa et al., 2011), TOPCAT (Taylor, 2005)
@@ 1 Introduction
Identifying WD+MS post-CE systems in stellar clusters can circumvent this issue by providing independent constraints on the age of the system. This, in turn, allows for more detailed probes of the pre-CE evolution. Such efforts have been successfully applied to two post-CE WD+MS binaries in the Hyades: V471 Tau and HZ9. For V471 Tau, Muirhead et al. (2022) estimate the pre-CE mass of the WD progenitor to be between $M=3.16M_{\odot}-3.41M_{\odot}$. Combined with the WD’s temperature and age of the Hyades, this progenitor mass implies that the system could not have formed via standard binary evolution. Both Muirhead et al. (2022) and O’Brien et al. (2001) favor a scenario where a merger occurred in a triple system and subsequently entered a CE with the tertiary star after the merged product evolved off the MS. In the case of HZ9, Muirhead et al. (2024) measure a current mass for the WD of $M=0.504\pm 0.016M_{\odot}$. Combined with the cooling age, this is compatible with isolated binary evolution for a system that underwent a CE phase.
@@
In this work, we perform the first systematic search for WD+MS binary systems in 299 Milky Way OCs. We define our OC and stellar samples in Section 2, and outline the supervised machine learning approach used to photometrically identify WD+MS binary candidates in this study in Section 3. Section 4 presents the photometric, spatial and kinematic properties of our high-probability WD+MS candidates, while follow-up spectra and archival light curves for an example subset of these systems are presented in Section 5. In Section 6.3, we discuss other possible astrophysical sources present in our catalogue by photometrically comparing our candidates to known samples of cataclysmic variables, MS+MS binaries, rapidly rotating (active) M-dwarfs and single WDs. We discuss implications of natal kicks in Section 6.4 and summarize in Section 7. Spectral energy distributions (SEDs) for all of our candidates are presented in Appendix C. Ultimately, this catalogue is an essential first step in identifying a population of WD+MS post-CE binaries in clusters.
@@ 2.1 Open Cluster Selection
Our sample of Milky Way OCs stems from the Cantat-Gaudin et al. (2020) catalogue – a compilation of 2017 OCs identified in Gaia DR2 (Gaia Collaboration et al., 2016, 2018a). Since we are searching for WD+MS systems, the OCs in our study must adhere to specific distance, declination and age criteria. First, we discard 150 OCs in this sample that do not have reported distances or ages. The rest of our selection criteria to produce our final OC sample is described below.
@@ 2.1.1 Distance
We place an upper limit on the distance to the clusters that we search in this study to ensure that (i) WD+MS binaries are detectable given the depth of the surveys we utilize and (ii) reliable Gaia parallaxes are expected to be broadly available. To assess the detectability of WD+MS binaries, we consider a known sample of spectroscopic WD+MS systems from the Sloan Digital Sky Survey (SDSS, Rebassa-Mansergas et al., 2010). While there are some bright outliers ($M_{G}\sim 6$ mag), these binaries generally have absolute magnitudes between $M_{G}=8-14$ mag in the Gaia G-band. At distances $\gtrsim 1.5$ kpc, bright WD+MS binaries would have magnitudes fainter than $m_{G}=19$ mag. Moreover, Gaia DR3 kinematics are essential for associating binaries with host clusters in this study, so it is imperative that distance measurements to the cluster stars are reliable. At large distances, using inverted parallax as a distance estimate is challenging due to the resulting skewed distance error distribution (Lutz & Kelker, 1973). Depending on colour and sky position, sources at distances further than $\sim 1.5$ kpc have errors in parallax $\gtrsim 20\%$ and are thus less reliable (Bailer-Jones, 2015; Bailer-Jones et al., 2021). Hence, we restrict ourselves to clusters at distances $<1.5$ kpc, as at distances greater than this, (i) even bright WD+MS binaries become harder to detect and (ii) parallax measurements become less reliable. Imposing this constraint on the full OC sample in Cantat-Gaudin et al. (2020) yields 589 clusters.
@@ 2.2.2 Kinematic Constraints
In this study, it is essential that the kinematics of each WD+MS candidate are consistent with those of its suspected host OC. While a complete kinematic analysis is applied to our full catalogue in Section 4.2, we initially apply the following set of rough kinematic constraints.
@@
We note that when searching for unresolved binaries in Gaia, it is often common to use constraints on the Renormalized Unit Weight Error (ruwe; e.g. Belokurov et al., 2020; Castro-Ginard et al., 2024). A ruwe of 1.0 is expected for single star systems, where ruwe values $>1.4$ indicate a single star model is a poor fit to the astrometric observations from Gaia (and hence, an unresolved multiple star system is possible). However, post-CE envelope binaries are believed to have extremely short orbital periods on the order of a few hours to a few days (Nebot Gómez-Morán et al., 2011). Small orbital separations would thus likely not be detected in the astrometric solution. As highlighted in El-Badry (2024), Gaia light curve data would be most powerful to identify short-period post-CE binaries, however, the sample of systems with light curve data is currently limited (but will be heavily expanded upon with the fourth data release of Gaia). While we do not consider ruwe when selecting stars for further analysis, we do provide ruwe values for each of the WD+MS candidates in our final catalogue (Section 4.3).
@@ 2.2.3 Photometric Constraints
    Gaia $BP-RP$ $\leq 2.6$: WD+MS binaries typically lie bluewards of the zero-age MS (ZAMS). This is especially evident in Rebassa-Mansergas et al. (2010), who present a catalogue of 1602 spectroscopically confirmed WD+MS systems with SDSS DR6 data. When cross-matching this catalogue with Gaia DR3, we find that 1468 systems have $BP-RP$ colour measurements, $>99.5\%$ of which have $BP-RP$ $<2.6$.
@@ 2.2.4 Cross-matching
For every OC in Section 2.1, we have a sample of Gaia DR3 stars from Section 2.2 that are kinematically similar to the host cluster and in the broad region of the Gaia colour-magnitude diagram (CMD) where WD+MS binaries are expected to appear. However, other classes of objects can also appear in this region of a Gaia CMD (see e.g. Section 6.3). Still, WD+MS systems exhibit specific ultraviolet, optical and infrared colours due to the presence of a hot WD and a cool M-dwarf companion (e.g. Rebassa-Mansergas et al., 2010). We therefore impose the added requirement that sufficient multi-band photometry is available to carry out a detailed assessment of their possible binary nature. To span a range of wavelengths, we cross-match each of our Gaia DR3 OC stellar samples with data from Pan-STARRS1 DR1 (Tonry et al., 2012; Chambers et al., 2016) and the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006) using TOPCAT (Taylor, 2005). We adopt the default cross-matching parameters in TOPCAT, using a cross-match radius of 1.0 arcsecond. We note that two clusters (UBC 577 and UPK 303) had zero cross-matches, so they are discarded from our sample. Finally, we only select stars that have greater than three-sigma detections ($\sigma_{mag}<0.36$) in all photometric bands to ensure the data is high quality.
While the inclusion of shorter-wavelength photometric bands (i.e. $u$-magnitudes) would allow for detection of hotter WDs (and would better separate WD+MS binaries from MS stars in colour-colour spaces), existing $u$-band data for OCs is limited. For instance, SDSS DR16 (Ahumada et al., 2020)—which contains $u$-band photometry—only offers sporadic coverage of the Galactic plane (where most OCs are located). Specifically, less than a few dozen of the 301 clusters that meet our criteria outlied in Section 2.1 retain $>50\%$ of their stars when cross-matching their stellar samples with SDSS DR16. Other ultraviolet/$u$-band photometric surveys like the Galaxy Evolution Explorer (GALEX; Martin et al., 2005) and the Ultraviolet Near Infrared Optical Northern Survey (UNIONS; Ibata et al., 2017) also primarily observe the Galactic halo. Future surveys like the Ultraviolet Explorer (UVEX; Kulkarni et al., 2021) will provide all sky ultraviolet coverage, allowing for incorporation of bluer wavelengths.
@@ 3 WD+MS Candidate Identification
As mentioned in Section 2.2.4, WD+MS binaries exhibit unique photometry and occupy unique regions of CMDs and colour-colour spaces. Previous studies that identified WD+MS binaries in the field (e.g. Willems & Kolb, 2004; Rebassa-Mansergas et al., 2010; Ren et al., 2014; Rebassa-Mansergas et al., 2021; Nayak et al., 2023) combined multi-wavelength observations in a variety of CMDs and colour-colour spaces, highlighting that these binaries are generally photometrically distinct from the majority of MS stars. Here, we present a supervised machine learning method to identify high-probability WD+MS binary candidates based on a variety of input colours. We adopt a machine learning as opposed to manual approach because, in providing probabilities, it can better handle edge cases and/or stars with individual photometric bands that are unphysical (e.g. due to catalogue errors). The machine learning method, training data, input features, model definition/evaluation, application to stellar samples and additional vetting of candidates are described below. The overall candidate identification procedure is summarized in Figure 1.
@@ 3.2.1 Cluster Star Training Sample
Figure 2: Properties of the training data used to identify WD+MS binaries. MS cluster members from five OCs (Alessi 10, ASCC 41, NGC 2251, NGC 2632 and NGC 7086) are indicated as gray circles. SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010) are marked as squares, where the sample is split into two classes. Class 1 (dark pink outlined) represents WD+MS binaries that are distinct from the MS cluster members in a t-SNE projection, five example colour-colour spaces and a Gaia CMD, while Class 2 (light purple) represent a subset of binaries that are more likely to share properties with MS cluster members. To maximize separation between the MS stars and WD+MS binaries and thus minimize potential contamination, we only use Class 1 WD+MS binaries when training our SVM.
@@ 3.2.2 WD+MS Training Sample
Our set of known WD+MS systems is taken from Rebassa-Mansergas et al. (2010), who provide a catalogue of 1602 WD+MS binaries from SDSS DR6. After cross-matching this catalogue with Gaia DR3, Pan-STARRS1 and 2MASS and cleaning the sample using the same constraints as the cluster star samples and training data, we are left with 570 WD+MS binaries.
To explore the global properties of the Rebassa-Mansergas et al. (2010) sample (which were identified spectroscopically using a template fitting technique), we use the t-SNE algorithm (van der Maaten & Hinton, 2008). T-SNE is an unsupervised machine learning method that reduces the dimensionality of a high-dimensional dataset and has been widely used in astronomy to easily examine and visualize similarities (e.g. Anders et al., 2018; Kos et al., 2018; Čotar et al., 2019; Grondin et al., 2023). Using scikit-learn (Pedregosa et al., 2011), we run the t-SNE algorithm with an input set of 10 colours from Gaia DR3, Pan-STARRS1 and 2MASS with a high perplexity ($p=100$), to identify global rather than local features. The results are plotted in Figure 2.
Figure 3: Our support vector machine (SVM) model to identify WD+MS binaries in five different colour-colour spaces. Our test data is run through the SVM, where each point is coloured by its probability of being a WD+MS binary. The training sample containing OC members (gray) and known WD+MS binaries from SDSS (purple, Rebassa-Mansergas et al., 2010) is highlighted for reference. Overall, the assigned probabilities for the test data are aligned with the SVM classifier (rainbow gradient), where any outliers can likely be explained by higher-dimensional effects (i.e. here, we use 2-D example spaces to visualize the SVM classification, however the SVM is actually a 10-dimensional hyperplane and uses all input colours to compute probabilities).
@@
The overall redder colours and higher absolute magnitudes of the Class 2 systems may simply be attributed to higher mass MS companions. However, after running a series of tests and multiple iterations of our SVM model, we chose to remove them from our training sample in order to minimize contamination and maximize separation between the cluster members and WD+MS binaries. We discuss the impact of this choice on our final sample contamination and completeness in Sections 3.5.1 and 3.5.2 below. Class 1 represents the final training sample of WD+MS binaries in this study.

### v2

# The first catalog of candidate white dwarf-main sequence binaries in open star clusters: A new window into common envelope evolution

@@ Abstract
Close binary systems are the progenitors to both Type Ia supernovae and the compact object mergers that can be detected via gravitational waves. To achieve a binary with a small radial separation, it is believed that the system likely undergoes common envelope (CE) evolution. Despite its importance, CE evolution may be one of the largest uncertainties in binary evolution due to a combination of computational challenges and a lack of observed benchmarks where both the post-CE and pre-CE conditions are known. Identifying post-CE systems in star clusters can partially circumvent this second issue by providing an independent age constraint on the system. For the first time, we conduct a systematic search for white dwarf (WD) and main-sequence (MS) binary systems in 299 Milky Way open star clusters. Coupling Gaia DR3 photometry and kinematics with multi-band photometry from Pan-STARRS1 and 2MASS, we apply a machine learning based approach and find 52 high-probability candidates in 38 open clusters. For a subset of our systems, we present follow-up spectroscopy from the Gemini and Lick Observatories and archival light curves from TESS, Kepler/K2 and the Zwicky Transient Facility. Examples of M-dwarfs with hot companions are spectroscopically observed, along with regular system variability. While the kinematics of our candidates are consistent with their host clusters, some systems have spatial positions offset relative to their hosts, potentially indicative of natal kicks. Ultimately, this catalog is a first step to obtaining a set of observational benchmarks to better link post-CE systems to their pre-CE progenitors.
@@
facilities: Gaia, Gemini (GMOS), Kepler/K2, Lick (Kast), Pan-STARRS, TESS, ZTF, 2MASSsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), lightkurve (Lightkurve Collaboration et al., 2018), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010), pyraf (Science Software Branch at STScI, 2012), scikit-learn (Pedregosa et al., 2011), TOPCAT (Taylor, 2005)
@@ 1 Introduction
Identifying WD+MS post-CE systems in stellar clusters can circumvent this issue by providing independent constraints on the age of the system. This, in turn, allows for more detailed probes of the pre-CE evolution. Such efforts have been successfully applied to two post-CE WD+MS binaries in the Hyades: V471 Tau and HZ9. For V471 Tau, Muirhead et al. (2022) estimate the pre-CE mass of the WD progenitor to be between $M=3.16M_{\odot}-3.41M_{\odot}$. Combined with the WD’s temperature and age of the Hyades, this progenitor mass implies that the system could not have formed via standard binary evolution. Both Muirhead et al. (2022) and O’Brien et al. (2001) favor a scenario where a merger occurred in a triple system and subsequently entered a CE with the tertiary star after the merged product evolved off the MS. In the case of HZ9, Muirhead et al., in prep measure a current mass for the WD of $M=0.504\pm 0.016M_{\odot}$. Combined with the cooling age, this is compatible with isolated binary evolution for a system that underwent a CE phase.
@@
In this work, we perform the first systematic search for WD+MS binary systems in 299 Milky Way OCs. We define our OC and stellar samples in Section 2, and outline the supervised machine learning approach used to photometrically identify WD+MS binary candidates in this study in Section 3. Section 4 presents the photometric, spatial and kinematic properties of our high-probability WD+MS candidates, while follow-up spectra and archival light curves for an example subset of these systems are presented in Section 5. In Section 6.3, we discuss other possible astrophysical sources present in our catalog by photometrically comparing our candidates to known samples of cataclysmic variables, MS+MS binaries, rapidly rotating (active) M-dwarfs and single WDs. We discuss implications of natal kicks in Section 6.4 and summarize in Section 7. Spectral energy distributions (SEDs) for all of our candidates are presented in Appendix C. Ultimately, this catalog is an essential first step in identifying a population of WD+MS post-CE binaries in clusters.
@@ 2.1 Open Cluster Selection
Our sample of Milky Way OCs stems from the Cantat-Gaudin et al. (2020) catalog – a compilation of 2017 OCs identified in Gaia DR2 (Gaia Collaboration et al., 2016, 2018a). Since we are searching for WD+MS systems, the OCs in our study must adhere to specific distance, declination and age criteria. First, we discard 150 OCs in this sample that do not have reported distances or ages. The rest of our selection criteria to produce our final OC sample is described below.
@@ 2.1.1 Distance
We place an upper limit on the distance to the clusters that we search in this study to ensure that (i) WD+MS binaries are detectable given the depth of the surveys we utilize and (ii) reliable Gaia parallaxes are expected to be broadly available. To assess the detectability of WD+MS binaries, we consider a known sample of spectroscopic WD+MS systems from the Sloan Digital Sky Survey (SDSS, Rebassa-Mansergas et al., 2010). While there are some bright outliers ($M_{G}\sim 6$ mag), these binaries generally have absolute magnitudes between $M_{G}=8-14$ mag in the Gaia G-band. At distances $\gtrsim 1.5$ kpc, bright WD+MS binaries would have magnitudes fainter than $m_{G}=19$ mag. Moreover, Gaia DR3 kinematics are essential for associating binaries with host clusters in this study, so it is imperative that distance measurements to the cluster stars are reliable. At large distances, using inverted parallax as a distance estimate is challenging due to the resulting skewed distance error distribution (Lutz & Kelker, 1973). Depending on color and sky position, sources at distances further than $\sim 1.5$ kpc have errors in parallax $\gtrsim 20\%$ and are thus less reliable (Bailer-Jones, 2015; Bailer-Jones et al., 2021). Hence, we restrict ourselves to clusters at distances $<1.5$ kpc, as at distances greater than this, (i) even bright WD+MS binaries become harder to detect and (ii) parallax measurements become less reliable. Imposing this constraint on the full OC sample in Cantat-Gaudin et al. (2020) yields 589 clusters.
@@ 2.2.2 Kinematic Constraints
In this study, it is essential that the kinematics of each WD+MS candidate are consistent with those of its suspected host OC. While a complete kinematic analysis is applied to our full catalog in Section 4.2, we initially apply the following set of rough kinematic constraints.
@@
We note that when searching for unresolved binaries in Gaia, it is often common to use constraints on the Renormalized Unit Weight Error (ruwe; e.g. Belokurov et al., 2020; Castro-Ginard et al., 2024). A ruwe of 1.0 is expected for single star systems, where ruwe values $>1.4$ indicate a single star model is a poor fit to the astrometric observations from Gaia (and hence, an unresolved multiple star system is possible). However, post-CE envelope binaries are believed to have extremely short orbital periods on the order of a few hours to a few days (Nebot Gómez-Morán et al., 2011). Small orbital separations would thus likely not be detected in the astrometric solution. As highlighted in El-Badry (2024), Gaia light curve data would be most powerful to identify short-period post-CE binaries, however, the sample of systems with light curve data is currently limited (but will be heavily expanded upon with the fourth data release of Gaia). While we do not consider ruwe when selecting stars for further analysis, we do provide ruwe values for each of the WD+MS candidates in our final catalog (Section 4.3).
@@ 2.2.3 Photometric Constraints
    Gaia $BP-RP$ $\leq 2.6$: WD+MS binaries typically lie bluewards of the zero-age MS (ZAMS). This is especially evident in Rebassa-Mansergas et al. (2010), who present a catalog of 1602 spectroscopically confirmed WD+MS systems with SDSS DR6 data. When cross-matching this catalog with Gaia DR3, we find that 1468 systems have $BP-RP$ color measurements, $>99.5\%$ of which have $BP-RP$ $<2.6$.
@@ 2.2.4 Cross-matching
For every OC in Section 2.1, we have a sample of Gaia DR3 stars from Section 2.2 that are kinematically similar to the host cluster and in the broad region of the Gaia color-magnitude diagram (CMD) where WD+MS binaries are expected to appear. However, other classes of objects can also appear in this region of a Gaia CMD (see e.g. Section 6.3). Still, WD+MS systems exhibit specific ultraviolet, optical and infrared colors due to the presence of a hot WD and a cool M-dwarf companion (e.g. Rebassa-Mansergas et al., 2010). We therefore impose the added requirement that sufficient multi-band photometry is available to carry out a detailed assessment of their possible binary nature. To span a range of wavelengths, we cross-match each of our Gaia DR3 OC stellar samples with data from Pan-STARRS1 DR1 (Tonry et al., 2012; Chambers et al., 2016) and the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006) using TOPCAT (Taylor, 2005). We adopt the default cross-matching parameters in TOPCAT, using a cross-match radius of 1.0 arcsecond. We note that two clusters (UBC 577 and UPK 303) had zero cross-matches, so they are discarded from our sample. Finally, we only select stars that have greater than three-sigma detections ($\sigma_{mag}<0.36$) in all photometric bands to ensure the data is high quality.
While the inclusion of shorter-wavelength photometric bands (i.e. $u$-magnitudes) would allow for detection of hotter WDs (and would better separate WD+MS binaries from MS stars in color-color spaces), existing $u$-band data for OCs is limited. For instance, SDSS DR16 (Ahumada et al., 2020)—which contains $u$-band photometry—only offers sporadic coverage of the Galactic plane (where most OCs are located). Specifically, less than a few dozen of the 301 clusters that meet our criteria outlied in Section 2.1 retain $>50\%$ of their stars when cross-matching their stellar samples with SDSS DR16. Other ultraviolet/$u$-band photometric surveys like the Galaxy Evolution Explorer (GALEX; Martin et al., 2005) and the Ultraviolet Near Infrared Optical Northern Survey (UNIONS; Ibata et al., 2017) also primarily observe the Galactic halo. Future surveys like the Ultraviolet Explorer (UVEX; Kulkarni et al., 2021) will provide all sky ultraviolet coverage, allowing for incorporation of bluer wavelengths.
@@ 3 WD+MS Candidate Identification
As mentioned in Section 2.2.4, WD+MS binaries exhibit unique photometry and occupy unique regions of CMDs and color-color spaces. Previous studies that identified WD+MS binaries in the field (e.g. Willems & Kolb, 2004; Rebassa-Mansergas et al., 2010; Ren et al., 2014; Rebassa-Mansergas et al., 2021; Nayak et al., 2023) combined multi-wavelength observations in a variety of CMDs and color-color spaces, highlighting that these binaries are generally photometrically distinct from the majority of MS stars. Here, we present a supervised machine learning method to identify high-probability WD+MS binary candidates based on a variety of input colors. We adopt a machine learning as opposed to manual approach because, in providing probabilities, it can better handle edge cases and/or stars with individual photometric bands that are unphysical (e.g. due to catalog errors). The machine learning method, training data, input features, model definition/evaluation, application to stellar samples and additional vetting of candidates are described below. The overall candidate identification procedure is summarized in Figure 1.
@@ 3.2.1 Cluster Star Training Sample
Figure 2: Properties of the training data used to identify WD+MS binaries. MS cluster members from five OCs (Alessi 10, ASCC 41, NGC 2251, NGC 2632 and NGC 7086) are indicated as gray circles. SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010) are marked as squares, where the sample is split into two classes. Class 1 (dark pink outlined) represents WD+MS binaries that are distinct from the MS cluster members in a t-SNE projection, five example color-color spaces and a Gaia CMD, while Class 2 (light purple) represent a subset of binaries that are more likely to share properties with MS cluster members. To maximize separation between the MS stars and WD+MS binaries and thus minimize potential contamination, we only use Class 1 WD+MS binaries when training our SVM.
@@ 3.2.2 WD+MS Training Sample
Our set of known WD+MS systems is taken from Rebassa-Mansergas et al. (2010), who provide a catalog of 1602 WD+MS binaries from SDSS DR6. After cross-matching this catalog with Gaia DR3, Pan-STARRS1 and 2MASS and cleaning the sample using the same constraints as the cluster star samples and training data, we are left with 570 WD+MS binaries.
To explore the global properties of the Rebassa-Mansergas et al. (2010) sample (which were identified spectroscopically using a template fitting technique), we use the t-SNE algorithm (van der Maaten & Hinton, 2008). T-SNE is an unsupervised machine learning method that reduces the dimensionality of a high-dimensional dataset and has been widely used in astronomy to easily examine and visualize similarities (e.g. Anders et al., 2018; Kos et al., 2018; Čotar et al., 2019; Grondin et al., 2023). Using scikit-learn (Pedregosa et al., 2011), we run the t-SNE algorithm with an input set of 10 colors from Gaia DR3, Pan-STARRS1 and 2MASS with a high perplexity ($p=100$), to identify global rather than local features. The results are plotted in Figure 2.
Figure 3: Our support vector machine (SVM) model to identify WD+MS binaries in five different color-color spaces. Our test data is run through the SVM, where each point is colored by its probability of being a WD+MS binary. The training sample containing OC members (gray) and known WD+MS binaries from SDSS (purple, Rebassa-Mansergas et al., 2010) is highlighted for reference. Overall, the assigned probabilities for the test data are aligned with the SVM classifier (rainbow gradient), where any outliers can likely be explained by higher-dimensional effects (i.e. here, we use 2-D example spaces to visualize the SVM classification, however the SVM is actually a 10-dimensional hyperplane and uses all input colors to compute probabilities).
@@
The overall redder colors and higher absolute magnitudes of the Class 2 systems may simply be attributed to higher mass MS companions. However, after running a series of tests and multiple iterations of our SVM model, we chose to remove them from our training sample in order to minimize contamination and maximize separation between the cluster members and WD+MS binaries. We discuss the impact of this choice on our final sample contamination and completeness in Sections 3.5.1 and 3.5.2 below. Class 1 represents the final training sample of WD+MS binaries in this study.